OpenStudy (anonymous):

any1 wanna fill the holes on the topic of reaction tables and Kc and Qc and all that jazz ?

OpenStudy (anonymous):

@Australopithecus if youre not too busy, and in the chatty mood, and not sleep deprived lol.

OpenStudy (anonymous):

so many medals Michael Phelps gonna hate.

OpenStudy (anonymous):

So i know Qc is like the "actual amount" and Kc is the amount at equilibrium. I know that we can convert these values from Qc-->Qp using PV=NRT and also the same for Kc-->Kp. I know we use reaction tables to determine how much of a product we can get from a given amount of reactant. (at this point if anything is wrong, please, correct me).Now when we cross into the realm of acids and bases, all I know is Kw=Ka*Kb.What i don't understand is how we get Ka, or Kb from a reaction table. Ka and Kb are a bit fuzzy for me in general.

OpenStudy (anonymous):

**jumping in the shower so no rush, and thanks in advanced*

OpenStudy (anonymous):

and i'm back

OpenStudy (australopithecus):

Ka and Kb are equilibrium constants for describing the strength an acid or base respectively \[K_a = \frac{[H_3O][A-]}{[HA]}\] For, HA + H2O ⇌ H3O + A- \[K_b = \frac{[OH-][BH+]}{B}\] For, B + H2O ⇌ OH- + BH+ The larger the Ka or Kb the stronger the acid or base, they are especially useful as you can convert them to pKa or pKb (by taking the -log of them) which values tell you the pH or pOH respectively at which 50% of a molecule is protonated and deportonted Concentrations are at equilibrium, so just use the values to figure them out using the formula above. If you are given pH or pOH you can use the following formulas to help you pKw = pH + pOH = 14 pH = -log[H3O+] pOH = -log[-OH] Kw = [H3O+][-OH]

OpenStudy (anonymous):

Ok cool, thank you for the info, and a lot of fog was just cleared so you secured your medal. When talking about the bases, the molecule has to have strictly OH in it for it to be considered a base, or can it have one of the other group 16 elements and st ill be considered a base?

OpenStudy (australopithecus):

In acid base chemistry there are Lewis acid and Lewis bases a Lewis acid is a molecule that accepts electrons a Lewis base is a molecule that donates electrons A sub category of that is Bronsted bases and acids a Bronsted acid is a molecule that donates hydrogens a Bronsted base is a molecule that donates hydroxides It does not have to be strictly group 16 to be a base, as you can see by the definition of a lewis acid and a lewis base. Right now you are dealing with bronsted acid and base theory. It is also important to know that Ka and Kb are both environmentally and temperature dependent. Also, that pH s based in an aqueous environment.

OpenStudy (anonymous):

thank you so much! I really appreciate it.

OpenStudy (australopithecus):

If you go further in chemistry you will see that pKa and pKb values are very useful in determining the reactivity of a chemical

OpenStudy (australopithecus):

at least in organic chem